OpenStudy (anonymous):

Suppose a friend tells you she paid a total of $16,368 for a car, and you'd like to know the car's list price (the price before taxes) so that you can compare prices at various dealers. Find the list price of the car if your friend bought the car in: a) Arizona, where the sales tax is 6.6% b) New York, where the sales tax is 8.25% c) a state where the sales tax is r d) solve for r in your answer to (c) above.

OpenStudy (jdoe0001):

well 16.368 is the (list price) + tax thus (list price) + (tax) = 16,368 so the list price will be 16,368 - tax what if the tax were 6.6%, or 8.25%? well \(\begin{array}{llll} \$&\% \\\hline\\ 16368&100\\ x&6.6 \end{array}\implies \cfrac{16368}{x}=\cfrac{100}{6.6}\implies x=? \\ \quad \\ \quad \\ \begin{array}{llll} \$&\% \\\hline\\ 16368&100\\ x&8.25 \end{array}\implies \cfrac{16368}{x}=\cfrac{100}{8.25}\implies x=? \\ \quad \\ list\ price=16368-x\)

OpenStudy (anonymous):

Thank you!!!